After a sailboat had been on the water for 3 h, a change in the wind direction reduced the average speed of the boat by 3 mph. The entire distance sailed was 54 mi. The total time spent sailing was 5 h. How far did the sailboat travel in the first 3 h
Answer by stanbon(75887)

After a sailboat had been on the water for 3 h, a change in the wind direction reduced the average speed of the boat by 3 mph. The entire distance sailed was 54 mi. The total time spent sailing was 5 h. How far did the sailboat travel in the first 3 h
----------------
1st leg DATA:
time 3 hr ; rate = x mph ; distance = 3x miles
----
2nd leg DATA:
time = 2 hr ; rate = x-3 mph ; distance = 2x-6 miles
-----
Equation:
dist + dist = 54 miles
3x + 2x - 6 = 54
5x = 60
x = 12 mph
---
Ans: distance traveled on 1st leg = 3x = 3*12 = 36 mph
--------------
Cheers,
Stan H.
--------------